A rod is inserted as the core in the current-carrying solenoid of the previous problem. (a) What is the magnetic intensity H at the centre? (b) If the magnetization I of the core is found to be 0.12 A/m, find the susceptibility of the material of the rod. (c) Is the material paramagnetic, diamagnetic or ferromagnetic?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. (a) H = 1500 A/m As the solenoid and the rod are long and we are interested in the magnetic intensity at the centre, the end effects may be neglected. There is no effect of the rod on the magnetic intensity at the centre (b) I = 0.12 A/m We know I ⃗ = XH ⃗ X = Susceptibility ⇒ X = I / H = 0.12/1500 = 0.00008 = 8x 10^-5 (c) The material is paramagnetic.